what is 6 5/8% as a fraction?
Answer by stanbon(75887)

what is 6 5/8% as a fraction?
----------
6 5/8% = [6 5/8]/100
---
= (53/8]/100
---
= 53/800
==============
Cheers,
Stan H.
==============